# HumanGif: Single-View Human Diffusion with Generative Prior

### v1

@@ Abstract.
While previous single-view-based 3D human reconstruction methods made significant progress in novel view synthesis, it remains a challenge to synthesize both view-consistent and pose-consistent results for animatable human avatars from a single image input. Motivated by the success of 2D character animation, we propose HumanGif, a single-view human diffusion model with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models. To ensure fine-grained and consistent novel view and pose synthesis, we introduce a Human NeRF module in HumanGif to learn spatially aligned features from the input image, implicitly capturing the relative camera and human pose transformation. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. Extensive experiments on RenderPeople and DNA-Rendering datasets demonstrate that HumanGif achieves the best perceptual performance, with better generalizability for novel view and pose synthesis. Code will be released at https://github.com/skhu101/HumanGif.
Diffusion Model, Generative Prior, Human NeRF
ccs: Computer systems organization Embedded systemsccs: Computer systems organization Redundancyccs: Computer systems organization Roboticsccs: Networks Network reliability
@@
Figure 1. HumanGif is a single-view human diffusion model. By inheriting generative prior, HumanGif synthesizes realistic view and pose-consistent images.
1. Introduction
@@
Synthesizing 3D human performers with consistent novel views and poses holds extensive utility across various domains, including AR/VR, video games, and movie production. Recent methods enable the novel view and pose synthesis of 3D human avatars from sparse-view human videos (Peng et al., 2021b; Chen et al., 2021; Peng et al., 2022; Xu et al., 2021; Noguchi et al., 2021; Weng et al., 2022a; Su et al., 2021; Jiang et al., 2022a, c; Wang et al., 2022), with Neural Radiance Field (Mildenhall et al., 2021) or Gaussian Splatting-based (Kerbl et al., 2023) representation. While novel view and pose synthesis from sparse-view videos achieve impressive results, generating view-consistent and pose-consistent outputs from a single image (Hu et al., 2023; Dong et al., 2023; Weng et al., 2023; Liao et al., 2023; Huang et al., 2023b; Pan et al., 2024; Wang et al., 2024; Ho et al., 2024; Chen et al., 2024; Cha et al., 2023; Casas and Comino-Trinidad, 2023) remains challenging due to the limited information available in this setting.
@@
To address the challenge of missing information from a single image, a line of research (Weng et al., 2023; Huang et al., 2023b; Pan et al., 2024; Casas and Comino-Trinidad, 2023) focuses on the novel view synthesis from a single human image by complementing human images or UV textures at unseen views (e.g., back view). To further learn animatable human avatars from a single image, another line of research (Hu et al., 2023; Shao et al., 2024) proposes synthesizing both novel views and novel poses from a single image through learning a generalizable HumanNeRF or diffusion models. However, such frameworks typically require multi-view human image collections as training datasets, and their performance is closely tied to the quality and scale of these datasets. While efforts to create high-quality multi-view human image datasets (Yu et al., 2021; Han et al., 2023; Cheng et al., 2023; Li et al., 2021; Cai et al., 2022; Peng et al., 2021c; Ho et al., 2023; Tang et al., 2023; Alldieck et al., 2018; Ionescu et al., 2013; Mehta et al., 2017; Peng et al., 2022; Habermann et al., 2020, 2021; Huang et al., 2024; Işık et al., 2023; Yang et al., 2023a) have accelerated in recent years, the number of human subjects in these datasets remains significantly smaller compared to multi-view object and scene image datasets (Deitke et al., 2023; Wu et al., 2023; Yu et al., 2023b; Tung et al., 2025; Dai et al., 2017; Zhou et al., 2018). This data disparity limits the generalization performance of single-image-based 3D human modeling frameworks.
@@
To mitigate this data-sparsity dilemma, recent research (Bhunia et al., 2023; Karras et al., 2023; Wang et al., 2023; Chang et al., 2023; Xu et al., 2024; Hu, 2024; Zhu et al., 2024; Zhang et al., 2024a; Kim et al., 2024; Liyuan et al., 2024) in 2D character animation involves generative prior (inherit pre-trained weights) from Text-to-image (T2I) diffusion models (e.g., Stable Diffusion (Rombach et al., 2021)) to assist the single-view based 2D human animation (novel pose synthesis). By inheriting the generative priors from diffusion models, these approaches demonstrate impressive generalization performance for novel human subjects and poses in 2D character animation tasks (e.g., dancing video generation) using only a modest amount of training data, such as hundreds of monocular dancing videos (Jafarian and Park, 2021). Motivated by these successes, it is desirable to explore the use of generative prior from T2I diffusion models in the single-view-based 3D human novel view/pose synthesis task.
@@
In this work, we propose HumanGif, a single-view conditioned human diffusion with generative prior. Specifically, we formulate the single-view-based 3D human novel view/pose synthesis as a single-view-conditioned human diffusion process. Yet, this promising avenue comes with two primary barriers: 1) How to integrate human information into the human diffusion framework? 2) How to ensure that the performance achieved in the latent space is equally effective in the image-level space for human diffusion models? In particular, one potential solution is to allow diffusion models to learn relative camera and human pose transformation information through cross-attention between the reference image and the target pose image. However, our experiments reveal that human diffusion models struggle with a) learning detailed information (e.g., logos on T-shirts) even when such details are present in the reference image, and b) achieving consistent view and pose results. To tackle the challenges, we draw inspiration from prior research (Hu et al., 2023; Wu et al., 2024) and introduce a Human NeRF module (Hu et al., 2023) to capture the explicit fine-grained features in the target space by transforming the human subject from the reference pose space to the target pose space using linear blend skinning (LBS) of SMPL (Loper et al., 2015) and integrating them with volume rendering. With the SMPL LBS transformation, the resulting rendered feature map provides a spatially aligned conditioning signal that encodes the relative camera and human pose transformation. To further enhance the view-consistency and pose-consistency, we incorporate the multi-view attention and temporal attention in our HumanGif. Additionally, we observe a discrepancy between the latent space and image space for human images, attributed to the Variational Autoencoder (VAE) (Kingma et al., 2019) used in diffusion models. Inspired by the Radiance field rendering loss employed in 3D object diffusion models (Müller et al., 2023), we construct an image-level loss by mapping the noise latent to the image space through a VAE decoder. This ensures consistent optimizations in both latent and image spaces. With the above modules incorporated, our HumanGif successfully learns detailed information from the reference image and achieves the best perceptual novel view/pose consistent performance. Our main contributions are as follows:
  1. (1)
@@
    We introduce HumanGif, a single-view-conditioned human diffusion model that effectively incorporates generative priors from foundational diffusion models.
  2. (2)
@@
    To ensure fine-grained and consistent novel view and pose synthesis, we integrate a Human NeRF module with multi-view attention and temporal attention.
  3. (3)
@@
    Extensive experiments on two human datasets demonstrate that HumanGif outperforms baseline methods in perceptual quality of novel view and pose synthesis results.
2. Related Work
@@
Diffusion Models. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have demonstrated remarkable performance in image synthesis tasks, such as text-to-image (T2I) generation (Balaji et al., 2022; Nichol et al., 2021; Saharia et al., 2022; Ramesh et al., 2022; Huang et al., 2023a; Rombach et al., 2022). To improve sample quality for conditional generation, classifier guidance (Dhariwal and Nichol, 2021) and classifier-free guidance (Ho and Salimans, 2022) are introduced by leveraging an explicitly trained classifier or score estimates from both conditional and unconditional generation. In this work, we propose utilizing the generative prior from latent diffusion models (e.g., Stable Diffusion (Rombach et al., 2022)) for novel view and pose synthesis of 3D humans from a single image.
@@
Novel View/Pose Synthesis from a Single Human Image. Although 2D human modelling (AlBahar et al., 2021; Lewis et al., 2021; Sarkar et al., 2021; Men et al., 2020; Fu et al., 2022; Jiang et al., 2022b, 2023b; Fu et al., 2023) has made substantial progress, it is still challenging to synthesizing 3D humans from monocular inputs, especially for a single human image input (Saito et al., 2019; Saito et al., 2020; He et al., 2020; Li et al., 2020b; Dong et al., 2022; Li et al., 2020a; Bozic et al., 2021; Yang et al., 2021; Bhatnagar et al., 2020a, b; Huang et al., 2020; He et al., 2021; zheng2021pamir; Xiu et al., 2022b, a; Alldieck et al., 2022; Corona et al., 2022). To implement the information missing from the single image input, these approaches (Weng et al., 2023; Huang et al., 2023b; Pan et al., 2024; Wang et al., 2024; Chen et al., 2024; Choi et al., 2022; Zhang et al., 2024b; Ho et al., 2024; Dong et al., 2023; Cha et al., 2023; Casas and Comino-Trinidad, 2023) typically leverage text-to-image (T2I) diffusion models, such as Stable Diffusion (Rombach et al., 2021), by employing subject-specific Score Distillation Sampling (SDS) (Poole et al., 2022) or fine-tuning a T2I diffusion model conditioned on a front-view image using multi-view human datasets or 3D scan datasets. This enables the synthesis of human images from unseen viewpoints, such as generating a back view from a given front view. To further synthesize animatable 3D human avatars, recent research (Hu et al., 2023; Shao et al., 2024) leverages the strong modeling capabilities of deep neural networks to learn generalizable features for Neural Radiance Field (NeRF) or diffusion models, enabling the synthesis of novel views and poses in a feed-forward manner. However, the scale of multi-view human datasets limits their generalization ability in novel view and pose synthesis. In this work, we propose to leverage the generative prior from diffusion models to enhance the synthesis of animatable 3D human avatars with a single image input.
@@
Diffusion Models for Character Animation. The task of animating 2D human images focuses on generating animation videos from one or more static human images (AlBahar et al., 2023; Cao et al., 2024; Chan et al., 2019; Fu et al., 2022; Jiang et al., 2023a; Prokudin et al., 2021; Ren et al., 2020; Sarkar et al., 2020; Siarohin et al., 2019, 2021; Yoon et al., 2021; Yu et al., 2023a; Zhang et al., 2022; Zhao and Zhang, 2022). With advancements in text-to-image diffusion models, recent studies (Bhunia et al., 2023; Karras et al., 2023; Wang et al., 2023; Chang et al., 2023; Xu et al., 2024; Hu, 2024; Zhu et al., 2024; Zhang et al., 2024a; Kim et al., 2024; Liyuan et al., 2024) have investigated the use of diffusion models for 2D character animation. These approaches typically use a reference human image and a sequence of target pose images (e.g., OpenPose (Cao et al., 2019), DWPose (Yang et al., 2023b), DensePose (Güler et al., 2018), or SMPL pose (Loper et al., 2015)) as conditioning inputs to pose encoders or ControlNet (Zhang et al., 2023), generating a sequence of target images aligned with the appearance of the reference human image and the poses of the target images. Motivated by the success of diffusion-based methods in 2D character animation, we adapt these methodologies for 3D/4D human modeling from a single image. Human4DiT (Shao et al., 2024) introduced a diffusion model that generates multi-view human animation videos from a single image, utilizing factorized image, view, and temporal modules, similar to our approach. However, the code for this model is not publicly available. We identified that our baseline multi-view video diffusion model could not capture fine-grained details and maintain multi-view consistency. We show that integrating generative prior and our Human NeRF module into the diffusion pipeline effectively resolves these challenges. Our concurrent work GAS (Lu et al., 2025) also utilizes diffusion models to enhance the novel view and pose synthesis from a Human NeRF module.
3. Our Approach
@@
Figure 2. HumanGif Framework. Given a single input human image and a target pose sequence, our HumanGif produces a sequence of target images aligned with the human subject in the input image and target human poses. To synthesize view-consistent and pose-consistent outputs, our HumanGif proposes to incorporate generative prior, a Human NeRF module, image-level loss, multi-view attention, and temporal attention.
3.1. Preliminary
Latent Diffusion Model (LDM) (Rombach et al., 2022) presents a novel class of diffusion models by integrating two distinct stochastic processes, i.e., diffusion and denoising, directly within the latent space. LDM learns a variational autoencoder (VAE) (Kingma, 2013; Van Den Oord et al., 2017) to establish the mapping from image space to latent space, reducing the diffusion model’s complexity. Specifically, give an image $\bm{x}$, the encoder maps the image to a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ and the decoder reconstructs it to image space $\bm{x}=\mathcal{D}(\bm{z}_{0})$. The diffusion process progressively adds Gaussian noise to the data $\bm{z}_{0}$ following a variance schedule $1-\alpha_{0},\dots,1-\alpha_{T}$ specified for different time steps, i.e.,
(1) $\bm{z}_{t}=\sqrt{\alpha_{t}}\bm{z}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}.$
@@ 3.1. Preliminary
The denoising process learns to denoise $\bm{z}_{t}$ to $\bm{z}_{t-1}$ by predicting the noise $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$ for each denoising step, where $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$ is the output from a neural network (e.g., UNet (Ronneberger et al., 2015)) and $\bm{c}$ denotes conditional signal, e.g., the text embedding from CLIP (Radford et al., 2021) model. The diffusion loss $\mathcal{L}_{\text{diff}}$ is constructed by calculating the expected mean squared error (MSE) over the actual noise $\bm{\epsilon}$ and predicted noise $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$, i.e.,
(2) $\mathcal{L}_{\text{diff-latent}}=\mathbb{E}_{\bm{z}_{t},\bm{\epsilon},t,\bm{c}}[w_{t}\Arrowvert\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}\Arrowvert^{2}],$
@@
SMPL (Loper et al., 2015) is a parametric human model, $M(\bm{\beta},\bm{\theta})$, where $\bm{\beta},\bm{\theta}$ control body shape and pose, respectively. In this work, we utilize the Linear Blend Skinning (LBS) algorithm employed in SMPL to transform points from canonical to posed spaces. For instance, a 3D point $\bm{p}^{c}$ in the canonical space is transformed into the posed space defined by pose $\bm{\theta}$ as $\bm{p}^{\text{tgt}}=\sum_{k=1}^{K}w_{k}(\bm{G}_{k}(\bm{J},\bm{\theta})\bm{p}^{c}+\bm{b}_{k}(\bm{J},\bm{\theta},\bm{\beta}))$, where $\bm{J}$ represents $K$ joint locations, $\bm{G}_{k}(\bm{J},\bm{\theta})$ is the transformation matrix of joint $k$, $\bm{b}_{k}(\bm{J},\bm{\theta},\bm{\beta})$ is the translation vector of joint $k$, and $w_{k}$ is the linear blend weight.
NeRF (Mildenhall et al., 2021) learns an implicit, continuous function that maps the 3D location $\bm{p}$ and unit direction $\bm{d}$ of a point to its volume density $\bm{\sigma}\in[0,\infty)$ and color value $\bm{c}\in[0,1]^{3}$, i.e., $F_{\Phi}:(\gamma(\bm{p}),\gamma(\bm{d}))\to(\bm{c},\bm{\sigma})$, where $F_{\Phi}$ is parameterized by a multi-layer perceptron (MLP) network, $\gamma$ denotes a predefined positional embedding applied to $\bm{p}$ and $\bm{d}$. To render the RGB color of pixels in the target view, rays are cast from the camera origin $\bm{o}$ through the pixel along the unit direction $\bm{d}$. Based on the classical volume rendering (Kajiya and Von Herzen, 1984), the expected color $\hat{C}(\bm{r})$ for a camera ray $\bm{r}(t)=\bm{o}+t\bm{d}$ is computed as
@@
3.2. Single-view conditioned Human Diffusion Model
Building on the success of leveraging generative priors from T2I diffusion models to enhance generalization for novel poses and human subjects in the 2D character animation (Hu, 2024), we propose utilizing generative priors for single-view 3D human novel view/pose synthesis task.
Denoising UNet. Our backbone, illustrated in Fig. 2, is a denoising U-Net that inherits both the architecture and pre-trained weights (generative prior) from Stable Diffusion (SD) 1.5 (Rombach et al., 2022). The vanilla SD UNet in Stable Diffusion comprises three main components: downsampling, middle, and upsampling blocks. Each block integrates multiple interleaved layers, including convolution layers for feature extraction, self-attention layers for spatial feature aggregation, and cross-attention layers that interact with CLIP text embeddings to guide the denoising process. To process multiple noise latents to produce human images that align with the subject in the given reference image and the specified target poses, we further incorporate the reference image and the target pose sequence through the following modules.
ReferenceNet. Building on recent advances in 2D character animation (Hu, 2024), as illustrated in Fig. 2, we integrate a copy of the SD denoising U-Net as the ReferenceNet to extract features from the reference image. Specifically, we replace the self-attention layer with a spatial-attention layer, enabling self-attention on the concatenated features from the denoising U-Net and ReferenceNet. Additionally, we incorporate a CLIP encoder to extract semantic features, which are fused with the features from the denoising U-Net using cross-attention modules.
Pose Encoder. There are various options for defining target poses, including OpenPose, DensePose, DWPose, and SMPL parametric poses. We follow (Zhu et al., 2024) to estimate SMPL mesh for the reference image, animate the SMPL mesh with the target SMPL pose parameters, and render normal images as the target guidance signal. The SMPL normal pose images are encoded through a Pose Encoder (Hu, 2024), which contains four convolution layers. The Pose Encoder output is added to the noise latent to provide view and pose guidance information.
Table 1. Quantitative comparison of our HumanGif and baseline methods on the RendePeople and DNA-Rendering datasets.

### v2

@@ Abstract
Previous 3D human creation methods have made significant progress in synthesizing view-consistent and temporally aligned results from sparse-view images or monocular videos. However, it remains challenging to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image, as limited information is available in the single-view input setting. Motivated by the success of 2D character animation, we propose HumanGif, a single-view human diffusion model with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. To ensure fine-grained and consistent novel view and pose synthesis, we introduce a Human NeRF module in HumanGif to learn spatially aligned features from the input image, implicitly capturing the relative camera and human pose transformation. Furthermore, we introduce an image-level loss during optimization to bridge the gap between latent and image spaces in diffusion models. Extensive experiments on RenderPeople and DNA-Rendering datasets demonstrate that HumanGif achieves the best perceptual performance, with better generalizability for novel view and pose synthesis. Our code will be available at https://github.com/skhu101/HumanGif.
@@
Figure 1: HumanGif is a single-view human diffusion model that leverages generative priors to compensate for information missing from the input image. With generative prior, HumanGif synthesizes perpetually realistic, view-consistent, and temporally coherent human avatars.
1 Introduction
@@
Synthesizing 3D human performers with consistent novel views and poses holds extensive utility across various domains, including AR/VR, video games, and movie production. Recent methods enable the novel view and pose synthesis of 3D human avatars from sparse-view human videos [88, 19, 90, 125, 85, 119, 109, 59, 61, 114, 50, 21, 45, 48, 46], with Neural Radiance Field [82] or Gaussian Splatting-based [66] representation. While impressive novel view and pose synthesis results can be achieved from sparse-view videos, generating perpetually realistic, view-consistent, and temporally coherent human avatars from a single image [49, 29, 120, 75, 53, 86, 116, 42, 20, 15, 14] remains challenging due to the limited information available in this setting.
@@
To address the challenge of complementing information missing from a single image, a line of research [120, 53, 86, 14] focuses on the novel view synthesis from a single human image by complementing human images or UV textures at unseen views (e.g., back view) through additional models (e.g., T2I diffusion models). To further learn view-consistent and temporally aligned human avatars from a single image, another line of research [49, 105] proposes synthesizing novel views and poses from a single image through learning a generalizable HumanNeRF or a human diffusion model from scratch. However, such frameworks normally require multi-view human image collections as training datasets, and their performance is closely tied to the quality and scale of these datasets. While efforts to create high-quality multi-view human image datasets [131, 38, 22, 73, 11, 89, 41, 110, 4, 56, 80, 90, 36, 37, 55, 57, 128] have accelerated in recent years, the number of human subjects in these datasets remains significantly smaller compared to multi-view object and scene image datasets [27, 122, 133, 111, 26, 142]. This data disparity limits the generalization performance of single-image-based 3D human modeling frameworks.
@@
To mitigate this data-sparsity dilemma, recent research [9, 65, 115, 18, 126, 47, 143, 137, 67, 76] in 2D character animation involves generative prior (inherit pre-trained weights) from Text-to-image (T2I) diffusion models (e.g., Stable Diffusion [96]) to assist the single-view based 2D human animation (novel pose synthesis). By inheriting the generative priors from diffusion models, these approaches demonstrate impressive generalization performance for novel human subjects and poses in 2D character animation tasks (e.g., dancing video generation) using only a modest amount of training data, such as hundreds of monocular dancing videos [58]. Motivated by these successes, it is desirable to explore the use of generative prior from T2I diffusion models in the single-view-based 3D human novel view/pose synthesis task.
@@
In this work, we propose HumanGif, a single-view conditioned human diffusion with generative prior. Specifically, we formulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information. Yet, this promising avenue comes with two primary barriers: 1) How to integrate the relative camera and human pose information into the human diffusion framework? 2) How to ensure that the performance achieved in the latent space is equally effective in the image-level space for human diffusion models? In particular, one potential solution is to allow diffusion models to learn relative camera and human pose transformation information through cross-attention between the reference image and the target pose image. However, our experiments reveal that human diffusion models struggle with a) learning detailed information (e.g., logos on T-shirts) even when such details are present in the reference image, and b) achieving view-consistent and temporally aligned results. To tackle the challenges, we draw inspiration from prior 3D reconstruction methods [49, 121] and introduce a Human NeRF module [49] to capture the explicit fine-grained features in the target space by transforming the human subject from the reference pose space to the target pose space using linear blend skinning (LBS) of SMPL [77] and integrating them with volume rendering. With the SMPL LBS transformation, the resulting rendered feature map provides a spatially aligned conditioning signal that encodes the relative camera and human pose transformation. To further enhance the view and temporal consistency, we incorporate the multi-view attention and temporal attention in our HumanGif. Additionally, we observe a discrepancy between the latent space and image space for human images, attributed to the Variational Autoencoder (VAE) [69] used in diffusion models. Inspired by the Radiance field rendering loss employed in 3D object diffusion models [83], we construct an image-level loss by mapping the noise latent to the image space through a VAE decoder. This ensures consistent optimizations in both latent and image spaces. With the above modules incorporated, our HumanGif successfully recovers fine-grained information from the reference image and achieves the best perceptually realistic, view-consistent, and temporally coherent results. Our main contributions are as follows:
  1. 1.
@@
    We introduce HumanGif, a single-view-conditioned human diffusion model that incorporates generative priors to compensate for information missing from the single input image.
  2. 2.
@@
    To learn perpetually realistic, view-consistent, and temporally aligned 3D human avatars, we integrate a Human NeRF module with multi-view attention and temporal attention.
  3. 3.
@@
    Extensive experiments on two human datasets demonstrate that HumanGif outperforms baseline methods in perceptual quality of novel view and pose synthesis.
2 Related Work
@@
Diffusion Models. Diffusion models [108, 44] have demonstrated remarkable performance in image synthesis tasks, especially for text-to-image (T2I) generation [6, 84, 99, 94, 51, 97]. To improve sample quality for conditional generation, classifier guidance [28] and classifier-free guidance [43] are introduced by leveraging an explicitly trained classifier or score estimates from both conditional and unconditional generation. In this work, we propose utilizing the generative prior from latent diffusion models (e.g., Stable Diffusion [97]) for novel view and pose synthesis of 3D humans from a single image.
@@
Novel View/Pose Synthesis from a Single Human Image. Although 2D human modelling [2, 71, 104, 81, 31, 62, 63, 32] has made substantial progress, it is still challenging to synthesizing 3D humans from monocular inputs, especially for a single human image input [100, 101, 39, 74, 30, 72, 10, 127, 7, 8, 54, 40, 141, 124, 123, 5, 25]. To complement the information missing from a single image input, these approaches [120, 53, 86, 116, 20, 23, 138, 42, 29, 15, 14] typically leverage text-to-image (T2I) diffusion models, such as Stable Diffusion [96], by employing subject-specific Score Distillation Sampling (SDS) [91] or fine-tuning a T2I diffusion model conditioned on a front-view image using multi-view human datasets or 3D scan datasets. This enables the synthesis of human images from unseen viewpoints, such as generating a back view from a given front view. To further learn view-consistent and temporally aligned human avatars from a single image, recent research [49, 105] leverages the strong modeling capabilities of deep neural networks to learn generalizable features for Neural Radiance Field (NeRF) or diffusion models, enabling the synthesis of novel views and poses in a feed-forward manner. However, the scale of multi-view human datasets limits their generalization ability in novel view and pose synthesis. In this work, we propose to utilize the generative prior from diffusion models to produce perpetually realistic, view-consistent, and temporally coherent human avatars from a single image.
@@
Diffusion Models for 2D Character Animation. 2D Character Animation aims at generating temporally coherent animation videos from one or more static human images [3, 12, 16, 31, 60, 92, 95, 103, 106, 107, 130, 132, 135, 140]. With advancements in text-to-image diffusion models, recent studies [9, 65, 115, 18, 126, 47, 143, 137, 67, 76] have investigated the use of diffusion models for 2D character animation. These approaches normally use a reference human image and a sequence of target pose images (e.g., OpenPose [13], DWPose [129], DensePose [34], or SMPL pose [77]) as conditional inputs to pose encoders or ControlNet [134], generating a sequence of target images aligned with the reference human image and target pose images. Motivated by the success of diffusion-based methods in 2D character animation, we adapt these methodologies for 3D/4D human modeling from a single image. Human4DiT [105] introduced a diffusion model that generates multi-view human animation videos from a single image, utilizing factorized image, view, and temporal modules, similar to our approach. However, the code for this model is not publicly available. We identified that our baseline multi-view video diffusion model could not capture fine-grained details and maintain multi-view consistency. We show that integrating generative prior and our Human NeRF module into the diffusion pipeline effectively resolves these challenges. Our concurrent work GAS [78] also utilizes diffusion models to enhance the novel view and pose synthesis performance from a Human NeRF module.
3 Our Approach
@@
Figure 2: HumanGif Framework. Given a single input human image and a target pose sequence, our HumanGif produces a sequence of target images aligned with the human subject in the input image and target human poses. To synthesize view-consistent and pose-consistent outputs, our HumanGif proposes to incorporate generative prior, a Human NeRF module, image-level loss, multi-view attention, and temporal attention.
3.1 Preliminary
Latent Diffusion Model (LDM) [97] presents a novel class of diffusion models by integrating two distinct stochastic processes, i.e., diffusion and denoising, directly within the latent space. LDM learns a variational autoencoder (VAE) [68, 113] to establish the mapping from image space to latent space, reducing the diffusion model’s complexity. Specifically, give an image $\bm{x}$, the encoder maps the image to a latent representation $\bm{z}_{0}=\mathcal{E}(\bm{x})$ and the decoder reconstructs it to image space $\bm{x}=\mathcal{D}(\bm{z}_{0})$. The diffusion process progressively adds Gaussian noise to the data $\bm{z}_{0}$ following a variance schedule $1-\alpha_{0},\dots,1-\alpha_{T}$ specified for different time steps, i.e.,
$\bm{z}_{t}=\sqrt{\alpha_{t}}\bm{z}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}.$ (1)
@@ 3.1 Preliminary
The denoising process learns to denoise $\bm{z}_{t}$ to $\bm{z}_{t-1}$ by predicting the noise $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$ for each denoising step, where $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$ is the output from a neural network (e.g., UNet [98]) and $\bm{c}$ denotes conditional signal, e.g., the text embedding from CLIP [93] model. The diffusion loss $\mathcal{L}_{\text{diff}}$ is constructed by calculating the expected mean squared error (MSE) over the actual noise $\bm{\epsilon}$ and predicted noise $\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})$, i.e.,
$\mathcal{L}_{\text{diff-latent}}=\mathbb{E}_{\bm{z}_{t},\bm{\epsilon},t,\bm{c}}[w_{t}\Arrowvert\bm{\epsilon}_{\Phi}(\bm{z}_{t},t,\bm{c})-\bm{\epsilon}\Arrowvert^{2}],$ (2)
@@
SMPL [77] is a parametric human model, $M(\bm{\beta},\bm{\theta})$, where $\bm{\beta},\bm{\theta}$ control body shape and pose, respectively. In this work, we utilize the Linear Blend Skinning (LBS) algorithm employed in SMPL to transform points from canonical to posed spaces. For instance, a 3D point $\bm{p}^{c}$ in the canonical space is transformed into the posed space defined by pose $\bm{\theta}$ as $\bm{p}^{\text{tgt}}=\sum_{k=1}^{K}w_{k}(\bm{G}_{k}(\bm{J},\bm{\theta})\bm{p}^{c}+\bm{b}_{k}(\bm{J},\bm{\theta},\bm{\beta}))$, where $\bm{J}$ represents $K$ joint locations, $\bm{G}_{k}(\bm{J},\bm{\theta})$ is the transformation matrix of joint $k$, $\bm{b}_{k}(\bm{J},\bm{\theta},\bm{\beta})$ is the translation vector of joint $k$, and $w_{k}$ is the linear blend weight.
NeRF [82] learns an implicit, continuous function that maps the 3D location $\bm{p}$ and unit direction $\bm{d}$ of a point to its volume density $\bm{\sigma}\in[0,\infty)$ and color value $\bm{c}\in[0,1]^{3}$, i.e., $F_{\Phi}:(\gamma(\bm{p}),\gamma(\bm{d}))\to(\bm{c},\bm{\sigma})$, where $F_{\Phi}$ is parameterized by a multi-layer perceptron (MLP) network, $\gamma$ denotes a predefined positional embedding applied to $\bm{p}$ and $\bm{d}$. To render the RGB color of pixels in the target view, rays are cast from the camera origin $\bm{o}$ through the pixel along the unit direction $\bm{d}$. Based on the classical volume rendering [64], the expected color $\hat{C}(\bm{r})$ for a camera ray $\bm{r}(t)=\bm{o}+t\bm{d}$ is computed as
@@
3.2 Single-view conditioned Human Diffusion Model
Motivated by the success of leveraging generative priors from T2I diffusion models in the 2D character animation [47], we propose utilizing generative priors for single-view-based 3D human novel view/pose synthesis tasks. Specifically, we reformulate the single-view-based 3D human novel view and pose synthesis as a single-view-conditioned human diffusion process, utilizing generative priors from foundational diffusion models to complement the missing information.
Denoising UNet. Our backbone, illustrated in Fig. 2, is a denoising U-Net that inherits both the architecture and pre-trained weights (generative prior) from Stable Diffusion (SD) 1.5 [97]. The vanilla SD UNet in Stable Diffusion comprises three main components: downsampling, middle, and upsampling blocks. Each block integrates multiple interleaved layers, including convolution layers for feature extraction, self-attention layers for spatial feature aggregation, and cross-attention layers that interact with CLIP text embeddings to guide the denoising process. To process multiple noise latents to produce human images that align with the subject in the given reference image and the specified target poses, we further incorporate the reference image and the target pose sequence through the following modules.
ReferenceNet. Building on recent advances in 2D character animation [47], as illustrated in Fig. 2, we integrate a copy of the SD denoising U-Net as the ReferenceNet to extract features from the reference image. Specifically, we replace the self-attention layer with a spatial-attention layer, enabling self-attention on the concatenated features from the denoising U-Net and ReferenceNet. Additionally, we incorporate a CLIP encoder to extract semantic features, which are fused with the features from the denoising U-Net using cross-attention modules.
Pose Encoder. There are various options for defining target poses, including OpenPose, DensePose, DWPose, and SMPL parametric poses. We follow [143] to estimate SMPL mesh for the reference image, animate the SMPL mesh with the target SMPL pose parameters, and render normal images as the target guidance signal. The SMPL normal pose images are encoded through a Pose Encoder [47], which contains four convolution layers. The Pose Encoder output is added to the noise latent to provide view and pose guidance information.
Table 1: Quantitative comparison of our HumanGif and baseline methods on the RendePeople and DNA-Rendering datasets.